### v7

# The Negation Of Singer’s Conjecture For The Sixth Algebraic Transfer

Abstract
Let $\mathscr{A}$ be the Steenrod algebra over the field of characteristic two, $\mathbb{F}_{2}.$ Denote by $GL(q)$ the general linear group of rank $q$ over $\mathbb{F}_{2}.$ The algebraic transfer, introduced by W. Singer [Math. Z. 202 (1989), 493-523], is a rather effective tool for unraveling the intricate structure of the (mod-2) cohomology of the Steenrod algebra, ${\rm Ext}_{\mathscr{A}}^{q,*}(\mathbb{F}_{2},\mathbb{F}_{2}).$ The Kameko homomorphism is one of the useful tools to study the dimension of the domain of the Singer transfer. Singer conjectured that the algebraic transfer is always a monomorphism. This conjecture has stood for nearly 40 years, and in this work we demonstrate that it fails in general. More precisely, by developing a novel algorithm in the computer algebra system OSCAR for computing $GL(q)$-invariants of the kernel of the Kameko homomorphism, we disprove Singer’s conjecture for bidegree $(6,6+36)$.
keywords:
Adams spectral sequences; Steenrod algebra; Hit problem; Algebraic transfer; OSCAR (computer algebra system)
2020 MSC:
55Q45, 55S10, 55S05, 55T15, 68W30
1 Introduction and statement of the main result
Introduction. Let $\mathbb{F}_{2}$ be the prime field with two elements. We use the shorthand $H^{*}(X)$ (resp. $H_{*}(X)$) for the singular cohomology (resp. homology) groups with coefficients in $\mathbb{F}_{2}$. The Steenrod algebra $\mathscr{A}$ is the algebra of all stable cohomology operations over $\mathbb{F}_{2}$ and plays a fundamental role in Algebraic Topology, particularly in stable homotopy theory. A central problem in this field is computing the stable homotopy groups of spheres. Despite many profound results, this problem remains challenging and is far from being fully solved. Researchers have developed deep theories and practical tools to understand and compute these groups. One of the most useful tools is the Adams spectral sequence, which approximates the 2-primary stable homotopy groups of the sphere spectrum $\mathbb{S}^{0}$. Its input is the cohomology of the Steenrod algebra, ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})=\bigoplus_{r\geq 0}{\rm Ext}^{q,r}_{\mathscr{A}}(\widetilde{H}^{*}(\mathbb{S}^{0})=\mathbb{F}_{2},\mathbb{F}_{2}),$ where $q$ is the homological degree and $r$ is the internal degree. For a deeper understanding of $\mathrm{Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}),$ readers may refer to papers such as [4, 5, 14, 15, 16, 19]. Within the scope of this paper, another efficient instrument that we are especially interested in is the Singer algebraic transfer, proposed by Singer in 1989 [31]. Before delving into the details of the Singer transfer, we will recall some pertinent aspects.
Let $\mathcal{V}^{q}$ denote a $q$-dimensional $\mathbb{F}_{2}$-vector space. Since $\mathbb{F}_{2}$ is a prime field of size two, $\mathcal{V}^{q}$ can be regarded as a rank-$q$ elementary abelian 2-group. It is well-known that $H^{*}(\mathcal{V}^{q})\cong S(\mathcal{V}_{*}^{q})$, the symmetric algebra of the dual space $\mathcal{V}_{*}^{q}\equiv H^{1}(\mathcal{V}^{q})$. We can choose $x_{1},x_{2},\ldots,x_{q}$ to be a basis of $H^{1}(\mathcal{V}^{q})$. In this case, $P_{q}:=H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}[x_{1},x_{2},\ldots,x_{q}]$, the connected $\mathbb{N}$-graded polynomial algebra on generators of degree $1,$ equipped with the canonical unstable algebra structure over $\mathscr{A}.$ By dualizing, the mod-2 homology $H_{*}(\mathcal{V}^{q})$ is a divided power algebra on $q$ generators. Let $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q})$ be the subspace of $H_{*}(\mathcal{V}^{q})$ consisting of all elements that are annihilated by all positive degree Steenrod operations. The group $GL(q)$ acts regularly on $\mathcal{V}^{q}$ and therefore on $P_{q}$ and $H_{*}(\mathcal{V}^{q}).$ This action commutes with that of the algebra $\mathscr{A}$ and so acts on $\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}$ and $\mathcal{P}_{\mathscr{A}}H_{*}(\mathcal{V}^{q}).$ With the idea that the structure of the Ext groups can be studied through modular invariant theory, Singer [31] formulated a homomorphism denoted as:
$\displaystyle Tr_{q}(\mathbb{F}_{2}=\widetilde{H}^{*}(\mathbb{S}^{0})):(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})))_{n}=$ $\displaystyle(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})\otimes\widetilde{H}_{*}(\mathbb{S}^{0})))_{n}$
$\displaystyle\longrightarrow{\rm Ext}^{q,q+n}_{\mathscr{A}}(\widetilde{H}^{*}(\mathbb{S}^{0}),\mathbb{F}_{2})={\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2}),$
Then, he proved that $Tr_{q}(\mathbb{F}_{2})$ is an isomorphism for $q=1,\,2,$ and that the ”total” transfer
$\mathsf{Tr_{*}}:\bigoplus_{q,\,n}(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})))_{n}\longrightarrow\bigoplus_{q,\,n}{\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$
forms a homomorphism of (bi-graded) algebras.
The domain of $Tr_{q}(\mathbb{F}_{2})$ is closely related to the structure of the tensor product $\mathbb{F}_{2}\otimes_{\mathscr{A}}H^{*}(\mathcal{V}^{q})\cong\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}.$ Indeed, we give $\mathbb{F}_{2}$ the trivial $\mathscr{A}$-module structure. That is, the unit in $\mathscr{A}$ acts as a unit, while $Sq^{k}(\mathbb{F}_{2})=0$ for any $k\geq 1.$ Let $\mathscr{A}^{>0}$ denote the positive degree part of $\mathscr{A},$ and put
$QP_{q}:=\mathbb{F}_{2}\otimes_{\mathscr{A}}P_{q}\cong\mathscr{A}/\mathscr{A}^{>0}\otimes_{\mathscr{A}}P_{q}\cong P_{q}/(\mathscr{A}^{>0}\cdot P_{q}),$
where $\mathscr{A}^{>0}\cdot P_{q}$ refers to the subspace of $P_{q}$ composed of all homogeneous polynomials of the form $\sum_{k\geq 1}Sq^{k}(f_{k}),$ with $Sq^{k}\in\mathscr{A}^{>0}$ and $f_{k}\in P_{q}.$ Note that
$\bigoplus_{n\geq 0}H^{n}(\mathcal{V}^{q})\cong\bigoplus_{n\geq 0}(P_{q})_{n}=P_{q},\ \ \bigoplus_{n\geq 0}(QP_{q})_{n}\cong QP_{q},$
where
$\begin{array}[]{ll}\vskip 6.0pt plus 2.0pt minus 2.0ptH^{n}(\mathcal{V}^{q})\cong(P_{q})_{n}=\big\langle\big\{f\in P_{q}:\ \mbox{$\displaystyle f$ is a homogeneous polynomial of degree $\displaystyle n$}\big\}\big\rangle,\\ (QP_{q})_{n}=\big\langle\big\{[f]\in QP_{q}:\ f\in(P_{q})_{n}\big\}\big\rangle.\end{array}$
In [28], we have showed that
$\dim(QP_{q})_{n}\;=\;\binom{n+q-1}{q-1}\;-\;\operatorname{rank}(M),$
where $M$ is the matrix whose columns are the coordinate vectors (relative to the monomial basis of $P_{q}$) of the degree-$n$ basis elements in $\mathscr{A}^{>0}\cdot P_{q}$. However, finding a closed formula for $\operatorname{rank}(M)$, and hence for $\dim(QP_{q})_{n}$, for general $q$ and $n$ appears to be infeasible at present. It is therefore crucial to find effective bounds on $\operatorname{rank}(M)$, which would yield bounds on $\dim(QP_{q})_{n}$ through the identity above. We refer the reader to our recent article [30], where this problem is treated and an explicit bound for $\dim(QP_{q})_{n}$ is derived.
Let now $[(QP_{q})_{n}]^{GL(q)}$ denote the subspace of $(QP_{q})_{n}$ comprising all $GL(q)$-invariants of degree $n.$ Consequently, the domain of the algebraic transfer is dual to the invariant $[(QP_{q})_{n}]^{GL(q)}$ for any $n.$ It should be noted that the bi-graded sum $\bigoplus_{q,n}[(QP_{q})_{n}]^{GL(q)}$ possesses a co-algebra structure. (This fact is derived from the co-algebra structure on $\bigoplus_{q}H^{*}(\mathcal{V}^{q}),$ which comes from the natural isomorphisms $H^{*}(\mathcal{V}^{q})\cong H^{*}(\mathcal{V}^{i})\otimes_{\mathbb{F}_{2}}H^{*}(\mathcal{V}^{j})$ with $i+j=q.$) Therefore, dualizing the co-algebra yields an algebraic structure on the domain of the total transfer $\mathsf{Tr}_{*}$, as previously mentioned.
Understanding the structure and computing the dimensions of $(QP_{q})_{n}$ and the invariant spaces $[(QP_{q})_{n}]^{GL(q)}$ are extremely difficult problems, if not impossible, even with modern computer algebra systems. The Peterson conjecture [20], which was proven by Wood [38], provides further insight into the graded vector space $QP_{q}$. This conjecture states that $QP_{q}$ is trivial in degrees $n$ if $\mu(n)>q$, where $\mu(n)$ denotes the minimal integer $\zeta$ for which $n$ can be written as $\sum_{1\leq j\leq\zeta}(2^{d_{j}}-1)$ for some positive integers $d_{j}$. In light of this result, we now focus on investigating the domain of the algebraic transfer when $\mu(n)\leq q$. Notably, the condition $\mu(n)\leq q$ is equivalent to the useful formulation $\alpha(n+q)\leq q$, where $\alpha(k)$ is the number of 1’s in the binary expansion of the integer $k$. This helps characterize the relevant “families” of $n$ that satisfy this condition.
Over the past nearly four decades, the Singer transfer and related aspects have been extensively studied by numerous authors (see, e.g., [1, 3, 7, 8, 9, 10, 11, 12, 13, 17, 18, 20, 21, 32, 33, 23, 26, 27, 28, 29, 30, 36, 37, 38]). In particular, in [1], Boardman showed that $Tr_{3}(\mathbb{F}_{2})$ is also an isomorphism. Remarkably, Singer [31] proved that the algebraic transfer fails to be surjective in bidegree $(5,14),$ after which he proposed the following ensuing unsettled conjecture.
Conjecture 1.1.
$Tr_{q}(\mathbb{F}_{2})$ is a one-to-one homomorphism for any $q.$
The conjecture is also very difficult to attack, partly because the calculation of both the domain and the codomain of $Tr_{q}(\mathbb{F}_{2})$ is not easy. The conjecture is known to hold for $q=1,\,2$ by Singer [31], for $q=3$ by Boardman [1], and for $q=4$ by our recent works [23, 24, 25]. In this paper, we demonstrate that it does not hold in general.
Statement of our main result. We refute Conjecture 1.1 in general by showing it fails in bidegree $(6,6+36)$. (Note that $\mu(36)=4<6$.) We proceed by explicitly computing the dimensions of the domain and codomain of $Tr_{6}(\mathbb{F}_{2})$ in internal degree $36$. The following results are obtained.
Theorem 1.2.
For $q=6$ and $n=36,$ we have
$\dim(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})))_{n}=2.$
According to Bruner [2], Chen [5] and Lin [15], we have
${\rm Ext}^{6,6+36}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})=\mathbb{F}_{2}\cdot t,\ \ t\neq 0.$
Combining this and Theorem 1.2, we get
Corollary 1.3.
Conjecture 1.1 does not hold for bidegree $(6,6+36).$
As $(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{q})))_{n}$ is dual to $[(QP_{q})_{n}]^{GL(q)},$ Theorem 1.2 is equivalent to the following technical theorem:
Theorem 1.4.
For $q=6$ and $n=36,$ we have
$(QP_{q})_{n}]^{GL(q)}=\mathbb{F}_{2}\cdot([\zeta_{1}],\,[\zeta_{2}]),$
where the polynomials $\zeta_{1}$ and $\zeta_{2}$ are determined as follows:
$\displaystyle\zeta_{1}$ $\displaystyle=x_{1}^{3}x_{2}^{5}x_{3}^{9}x_{4}^{16}x_{5}x_{6}^{2}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{24}x_{5}x_{6}^{2}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{24}x_{5}x_{6}^{2}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{24}x_{5}x_{6}^{2}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}^{28}x_{5}x_{6}^{2}+x_{1}x_{2}^{3}x_{3}x_{4}^{28}x_{5}x_{6}^{2}+x_{1}x_{2}x_{3}^{3}x_{4}^{28}x_{5}x_{6}^{2}+x_{1}x_{2}x_{3}x_{4}^{30}x_{5}x_{6}^{2}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{9}x_{4}x_{5}^{16}x_{6}^{2}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}x_{5}^{24}x_{6}^{2}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}x_{5}^{24}x_{6}^{2}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}x_{5}^{24}x_{6}^{2}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}x_{5}^{28}x_{6}^{2}+x_{1}x_{2}^{3}x_{3}x_{4}x_{5}^{28}x_{6}^{2}+x_{1}x_{2}x_{3}^{3}x_{4}x_{5}^{28}x_{6}^{2}+x_{1}x_{2}x_{3}x_{4}x_{5}^{30}x_{6}^{2}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{10}x_{4}x_{5}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{8}x_{5}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}x_{6}^{24}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}x_{6}^{24}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{2}x_{5}x_{6}^{24}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{4}x_{5}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{4}x_{5}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{4}x_{5}x_{6}^{24}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}x_{5}^{2}x_{6}^{24}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}x_{4}x_{5}^{4}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}x_{5}^{4}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}x_{5}^{4}x_{6}^{24}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}x_{5}x_{6}^{26}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}x_{4}x_{5}x_{6}^{26}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}x_{6}^{26}+x_{1}x_{2}x_{3}^{6}x_{4}x_{5}x_{6}^{26}+x_{1}^{3}x_{2}x_{3}x_{4}^{4}x_{5}x_{6}^{26}$
$\displaystyle\quad+x_{1}x_{2}x_{3}x_{4}^{6}x_{5}x_{6}^{26}+x_{1}^{3}x_{2}x_{3}x_{4}x_{5}^{4}x_{6}^{26}+x_{1}x_{2}x_{3}x_{4}x_{5}^{6}x_{6}^{26},$
$\displaystyle\zeta_{2}$ $\displaystyle=x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}x_{5}^{12}x_{6}^{4}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{13}x_{5}^{12}x_{6}^{4}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{12}x_{5}^{13}x_{6}^{4}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{14}x_{5}^{3}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{6}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{10}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{14}x_{5}^{6}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{4}x_{5}^{8}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{12}x_{5}^{8}x_{6}^{5}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{12}x_{5}^{8}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{12}x_{5}^{8}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{5}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{10}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{8}x_{5}^{12}x_{6}^{5}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{8}x_{5}^{12}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{8}x_{5}^{12}x_{6}^{5}+x_{1}^{7}x_{2}x_{3}x_{4}^{10}x_{5}^{12}x_{6}^{5}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}x_{4}^{10}x_{5}^{12}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{13}x_{6}^{5}+x_{1}x_{2}x_{3}^{6}x_{4}^{10}x_{5}^{13}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{14}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{14}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{5}+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{10}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{5}x_{6}^{6}+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{5}x_{5}^{6}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{9}x_{5}^{6}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{9}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{9}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{13}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{9}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{9}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{9}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{5}x_{5}^{12}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{13}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{13}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{7}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{12}x_{6}^{7}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{12}x_{6}^{7}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{14}x_{6}^{7}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{14}x_{6}^{7}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{7}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{5}x_{3}^{9}x_{4}^{9}x_{5}^{2}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{11}x_{5}^{3}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{14}x_{5}^{3}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{13}x_{5}^{4}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{13}x_{5}^{4}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{4}x_{5}^{5}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{12}x_{5}^{5}x_{6}^{8}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{12}x_{5}^{5}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{11}x_{5}^{6}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{13}x_{5}^{6}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}^{9}x_{4}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{11}x_{4}x_{5}^{8}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{5}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{5}x_{5}^{8}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{9}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}x_{4}^{9}x_{5}^{8}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{9}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{9}x_{5}^{8}x_{6}^{8}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{9}x_{5}^{8}x_{6}^{8}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{9}x_{5}^{8}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{8}x_{5}^{9}x_{6}^{8}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{9}x_{6}^{8}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{9}x_{5}^{9}x_{6}^{8}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{9}x_{5}^{9}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{9}x_{6}^{8}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{11}x_{5}^{9}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{11}x_{5}^{9}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{9}x_{4}x_{5}^{10}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{10}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{5}x_{5}^{10}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{9}x_{5}^{10}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{8}x_{5}^{11}x_{6}^{8}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{9}x_{5}^{11}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{9}x_{5}^{11}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{5}x_{5}^{12}x_{6}^{8}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{5}x_{5}^{12}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{4}x_{5}^{13}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{5}x_{5}^{13}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{5}x_{5}^{13}x_{6}^{8}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{13}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{7}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{10}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}^{7}x_{3}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{8}x_{5}^{8}x_{6}^{9}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{8}x_{5}^{8}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{8}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{9}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{9}x_{5}^{8}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{9}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{10}x_{5}^{8}x_{6}^{9}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{11}x_{5}^{8}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{11}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{8}x_{5}^{9}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{9}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{8}x_{5}^{9}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{8}x_{5}^{9}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{10}x_{5}^{9}x_{6}^{9}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{10}x_{5}^{9}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{10}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{9}x_{5}^{10}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{9}x_{5}^{10}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{14}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{14}x_{6}^{9}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{14}x_{6}^{9}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{14}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{8}x_{3}^{3}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{14}x_{3}^{3}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{8}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{8}x_{4}^{5}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}^{10}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{11}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{6}x_{3}^{11}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{14}x_{4}^{5}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{8}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{8}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{7}x_{4}^{9}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}^{7}x_{4}^{9}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{12}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{12}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{13}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{13}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{13}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{13}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{14}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{11}x_{5}^{4}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{13}x_{5}^{4}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{8}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{10}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{11}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{11}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{9}x_{5}^{6}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{11}x_{5}^{6}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{11}x_{5}^{6}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{13}x_{5}^{6}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{13}x_{5}^{6}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{13}x_{5}^{6}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{7}x_{5}^{8}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{7}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{9}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{9}x_{5}^{8}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{4}x_{4}^{3}x_{5}^{9}x_{6}^{10}+x_{1}^{7}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{7}x_{4}^{3}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}^{7}x_{4}^{3}x_{5}^{9}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{4}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{4}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{4}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{6}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{7}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{7}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{7}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{7}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{8}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{5}x_{5}^{10}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{11}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{10}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{7}x_{5}^{12}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{13}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{13}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{8}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{8}x_{5}^{8}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{8}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{9}x_{5}^{8}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{9}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{10}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{10}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{12}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}^{14}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}^{14}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{2}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}x_{4}^{6}x_{5}^{14}x_{6}^{11}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{14}x_{6}^{11}+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{4}x_{5}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{8}x_{4}^{5}x_{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{8}x_{4}^{5}x_{5}x_{6}^{12}+x_{1}^{7}x_{2}x_{3}^{10}x_{4}^{5}x_{5}x_{6}^{12}+x_{1}x_{2}^{7}x_{3}^{10}x_{4}^{5}x_{5}x_{6}^{12}+x_{1}^{3}x_{2}^{4}x_{3}^{11}x_{4}^{5}x_{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}^{11}x_{4}^{5}x_{5}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{8}x_{5}x_{6}^{12}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{12}x_{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{12}x_{5}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{12}x_{5}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{12}x_{5}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{13}x_{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{13}x_{5}x_{6}^{12}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{13}x_{5}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{13}x_{5}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{7}x_{5}^{3}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{8}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{11}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{11}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{11}x_{5}^{3}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{11}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{14}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{14}x_{5}^{3}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{14}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}x_{5}^{4}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{9}x_{5}^{4}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{9}x_{5}^{4}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{9}x_{5}^{4}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{11}x_{5}^{4}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{13}x_{5}^{4}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{12}x_{4}x_{5}^{5}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{8}x_{5}^{5}x_{6}^{12}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{8}x_{5}^{5}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{8}x_{5}^{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{9}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{9}x_{5}^{5}x_{6}^{12}+x_{1}^{7}x_{2}x_{3}x_{4}^{10}x_{5}^{5}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{10}x_{5}^{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}x_{4}^{10}x_{5}^{5}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}x_{3}^{7}x_{4}^{10}x_{5}^{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{11}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}^{6}x_{3}x_{4}^{11}x_{5}^{5}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{11}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{12}x_{5}^{5}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{12}x_{5}^{5}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{6}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{7}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{7}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{7}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{3}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{3}x_{5}^{8}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{5}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{5}x_{5}^{8}x_{6}^{12}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{5}x_{5}^{8}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{5}x_{5}^{8}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{5}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{9}x_{6}^{12}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{9}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{9}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{9}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{10}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{7}x_{5}^{10}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{7}x_{5}^{10}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{7}x_{5}^{10}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{11}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}x_{5}^{12}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}x_{5}^{12}x_{6}^{12}+x_{1}^{7}x_{2}x_{3}x_{4}^{3}x_{5}^{12}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{3}x_{5}^{12}x_{6}^{12}+x_{1}x_{2}^{7}x_{3}x_{4}^{3}x_{5}^{12}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{12}+x_{1}x_{2}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{5}x_{5}^{12}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{13}x_{6}^{12}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{3}x_{5}^{13}x_{6}^{12}+x_{1}x_{2}^{6}x_{3}x_{4}^{3}x_{5}^{13}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{3}x_{5}^{13}x_{6}^{12}+x_{1}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{13}x_{6}^{12}+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{13}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{12}x_{5}^{4}x_{6}^{13}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{5}x_{6}^{13}+x_{1}x_{2}x_{3}^{6}x_{4}^{10}x_{5}^{5}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{6}x_{6}^{13}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{13}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{4}x_{5}^{8}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{5}x_{5}^{8}x_{6}^{13}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{5}x_{5}^{8}x_{6}^{13}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{8}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{10}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{10}x_{6}^{13}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{10}x_{6}^{13}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{10}x_{6}^{13}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{12}x_{6}^{13}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{3}x_{5}^{12}x_{6}^{13}$
$\displaystyle\quad+x_{1}x_{2}^{6}x_{3}x_{4}^{3}x_{5}^{12}x_{6}^{13}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{3}x_{5}^{12}x_{6}^{13}+x_{1}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{13}+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{12}x_{6}^{13}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{12}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{5}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{5}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{7}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{14}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{9}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{9}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{14}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}^{11}x_{6}^{14}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{2}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{4}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{4}x_{5}^{11}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}^{6}x_{5}^{11}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{11}x_{6}^{14}+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{11}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}x_{5}^{14}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}x_{5}^{14}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}x_{5}^{14}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}x_{4}^{3}x_{5}^{14}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}x_{4}^{3}x_{5}^{14}x_{6}^{14}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{3}x_{4}^{3}x_{5}^{14}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{9}x_{5}x_{6}^{16}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{9}x_{5}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{9}x_{4}x_{5}^{2}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{9}x_{5}^{2}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{9}x_{5}^{2}x_{6}^{16}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{9}x_{5}^{2}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{8}x_{5}^{3}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{3}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{9}x_{5}^{3}x_{6}^{16}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}^{7}x_{3}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}x_{5}^{8}x_{6}^{16}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{7}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{3}x_{5}^{8}x_{6}^{16}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{5}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}x_{4}^{7}x_{5}^{8}x_{6}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}^{9}x_{6}^{16}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{9}x_{6}^{16}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}^{9}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{2}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{2}x_{5}^{9}x_{6}^{16}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{7}x_{4}^{2}x_{5}^{9}x_{6}^{16}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{3}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}^{6}x_{3}x_{4}^{3}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{16}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}^{6}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{9}x_{6}^{16}+x_{1}x_{2}x_{3}^{2}x_{4}^{7}x_{5}^{9}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}^{11}x_{6}^{16}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{11}x_{6}^{16}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{8}x_{5}^{3}x_{6}^{17}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{3}x_{6}^{17}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}^{8}x_{6}^{17}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}^{8}x_{6}^{17}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}^{8}x_{6}^{17}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{2}x_{5}^{8}x_{6}^{17}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{2}x_{5}^{8}x_{6}^{17}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{7}x_{4}^{2}x_{5}^{8}x_{6}^{17}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{3}x_{5}^{8}x_{6}^{17}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{8}x_{6}^{17}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{4}x_{5}^{8}x_{6}^{17}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{8}x_{6}^{17}+x_{1}x_{2}x_{3}^{2}x_{4}^{7}x_{5}^{8}x_{6}^{17}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{2}x_{5}^{9}x_{6}^{17}+x_{1}x_{2}^{6}x_{3}x_{4}^{2}x_{5}^{9}x_{6}^{17}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{2}x_{5}^{9}x_{6}^{17}+x_{1}x_{2}x_{3}^{6}x_{4}^{2}x_{5}^{9}x_{6}^{17}+x_{1}x_{2}x_{3}^{2}x_{4}^{6}x_{5}^{9}x_{6}^{17}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}^{10}x_{6}^{17}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{10}x_{6}^{17}+x_{1}x_{2}x_{3}^{2}x_{4}^{5}x_{5}^{10}x_{6}^{17}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}x_{5}^{8}x_{6}^{18}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}x_{5}^{8}x_{6}^{18}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}x_{5}^{8}x_{6}^{18}+x_{1}x_{2}x_{3}^{3}x_{4}^{5}x_{5}^{8}x_{6}^{18}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}^{9}x_{6}^{18}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{9}x_{6}^{18}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}^{4}x_{5}^{9}x_{6}^{18}+x_{1}x_{2}^{3}x_{3}x_{4}^{4}x_{5}^{9}x_{6}^{18}+x_{1}x_{2}x_{3}^{2}x_{4}^{5}x_{5}^{9}x_{6}^{18}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}^{8}x_{6}^{19}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{4}x_{4}x_{5}^{8}x_{6}^{19}+x_{1}x_{2}x_{3}^{7}x_{4}^{2}x_{5}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{4}x_{5}x_{6}^{24}+x_{1}^{7}x_{2}x_{3}x_{4}x_{5}^{2}x_{6}^{24}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}x_{4}x_{5}^{2}x_{6}^{24}+x_{1}x_{2}^{7}x_{3}x_{4}x_{5}^{2}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}x_{5}^{2}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}x_{5}^{2}x_{6}^{24}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{7}x_{4}x_{5}^{2}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}x_{4}^{5}x_{5}^{2}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}x_{4}^{5}x_{5}^{2}x_{6}^{24}+x_{1}x_{2}x_{3}^{3}x_{4}^{5}x_{5}^{2}x_{6}^{24}$
$\displaystyle\quad+x_{1}x_{2}x_{3}x_{4}^{7}x_{5}^{2}x_{6}^{24}+x_{1}x_{2}x_{3}^{2}x_{4}^{5}x_{5}^{3}x_{6}^{24}+x_{1}x_{2}x_{3}x_{4}^{6}x_{5}^{3}x_{6}^{24}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}x_{5}^{4}x_{6}^{24}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{3}x_{4}x_{5}^{4}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}x_{5}^{4}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}x_{4}^{3}x_{5}^{4}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}x_{4}^{2}x_{5}^{5}x_{6}^{24}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{2}x_{4}^{3}x_{5}^{5}x_{6}^{24}+x_{1}^{3}x_{2}x_{3}x_{4}x_{5}^{6}x_{6}^{24}+x_{1}x_{2}^{3}x_{3}x_{4}x_{5}^{6}x_{6}^{24}+x_{1}x_{2}x_{3}^{3}x_{4}x_{5}^{6}x_{6}^{24}$
$\displaystyle\quad+x_{1}x_{2}x_{3}x_{4}^{3}x_{5}^{6}x_{6}^{24}+x_{1}x_{2}x_{3}x_{4}^{2}x_{5}^{7}x_{6}^{24}+x_{1}x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{3}x_{6}^{25}+x_{1}x_{2}^{3}x_{3}x_{4}^{2}x_{5}^{4}x_{6}^{25}$
$\displaystyle\quad+x_{1}x_{2}x_{3}x_{4}^{2}x_{5}^{6}x_{6}^{25}+x_{1}x_{2}x_{3}x_{4}^{3}x_{5}^{4}x_{6}^{26}+x_{1}x_{2}x_{3}x_{4}x_{5}^{6}x_{6}^{26}+x_{1}x_{2}x_{3}x_{4}^{2}x_{5}^{4}x_{6}^{27}$
$\displaystyle\quad+x_{1}x_{2}x_{3}^{3}x_{4}^{2}x_{5}x_{6}^{28}+x_{1}x_{2}x_{3}x_{4}^{2}x_{5}^{3}x_{6}^{28}+x_{1}x_{2}x_{3}x_{4}x_{5}^{2}x_{6}^{30}.$
Remark 1.5.
To prove Theorem 1.4, we construct and implement a new algorithm in the OSCAR computer algebra system [39]. The algorithm computes an explicit basis for both the kernel of the Kameko homomorphism and the space $(QP_{q})_{n},$ as well as their corresponding invariants, for any $q$ and $n$ where $n-q$ is even. (Our previous algorithm in [29], implemented in SageMath, did not perform these basis and invariant computations for the kernel of the Kameko homomorphism.) Our reasoning for choosing OSCAR over SageMath for this implementation is detailed in Note 3.5(C) of Section 3. Furthermore, we used this new algorithm to verify previously known results, including those we computed by hand and those published by other authors (see, e.g., [3, 10, 26, 33, 35]). Our algorithm’s output is consistent with these established findings.
For instance, let us consider the case $q=5,\,n=35.$ In [33], Nguyen Sum had only determined the dimension of the invariant space $[(QP_{5})_{35}]^{GL(5)},$ to be one, without providing an explicit basis. This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [21, 22]). Such manual computations have since become obsolete. Our algorithm not only confirms this dimension but also provides an explicit basis for this space, including the dimension and basis of the invariant subspace of the kernel of the Kameko homomorphism. In particular, for $q=5$ and degree $35,$ our algorithm finds that the invariant space of the kernel of the Kameko homomorphism is trivial, while the invariant space $[(QP_{5})_{35}]^{GL(5)}$ is one-dimensional. The algorithm further shows that $[(QP_{5})_{35}]^{GL(5)}=\mathbb{F}_{2}\cdot[\texttt{GL5[1]}],$ where
GL5[1] $\displaystyle=\psi(q)+x_{1}^{15}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{6}+x_{1}^{7}x_{2}^{11}x_{3}^{5}x_{4}^{6}x_{5}^{6}+x_{1}^{3}x_{2}^{15}x_{3}^{5}x_{4}^{6}x_{5}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{6}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}^{7}x_{3}^{13}x_{4}^{6}x_{5}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{15}x_{4}^{6}x_{5}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{14}x_{5}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{14}x_{5}^{6}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{15}x_{5}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{10}x_{5}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{14}x_{5}^{10}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{7}x_{5}^{14}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{6}x_{5}^{15}+x_{1}^{7}x_{2}^{3}x_{3}x_{4}^{8}x_{5}^{16}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{8}x_{5}^{16}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{16}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}x_{3}^{7}x_{4}^{8}x_{5}^{16}+x_{1}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{16}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{9}x_{5}^{16}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{8}x_{5}^{17}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{17}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{8}x_{5}^{17}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{8}x_{5}^{17}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{10}x_{5}^{17}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{10}x_{5}^{17}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{8}x_{5}^{19}+x_{1}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{19}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}x_{5}^{24}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}^{6}x_{3}^{3}x_{4}x_{5}^{24}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}x_{5}^{24}+x_{1}^{7}x_{2}x_{3}x_{4}^{2}x_{5}^{24}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{2}x_{5}^{24}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{2}x_{5}^{24}+x_{1}x_{2}x_{3}^{7}x_{4}^{2}x_{5}^{24}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{3}x_{5}^{24}+x_{1}x_{2}^{6}x_{3}x_{4}^{3}x_{5}^{24}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{3}x_{5}^{24}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{4}x_{5}^{24}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{4}x_{5}^{24}+x_{1}^{3}x_{2}x_{3}^{2}x_{4}^{5}x_{5}^{24}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}^{3}x_{3}^{2}x_{4}^{5}x_{5}^{24}+x_{1}x_{2}^{3}x_{3}x_{4}^{6}x_{5}^{24}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}^{2}x_{5}^{25}+x_{1}x_{2}^{6}x_{3}x_{4}^{2}x_{5}^{25}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}x_{3}^{4}x_{4}^{2}x_{5}^{25}+x_{1}x_{2}x_{3}^{6}x_{4}^{2}x_{5}^{25}+x_{1}x_{2}x_{3}^{2}x_{4}^{6}x_{5}^{25}+x_{1}^{3}x_{2}^{4}x_{3}x_{4}x_{5}^{26}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}^{6}x_{3}x_{4}x_{5}^{26}+x_{1}^{3}x_{2}x_{3}^{4}x_{4}x_{5}^{26}+x_{1}x_{2}^{3}x_{3}x_{4}^{4}x_{5}^{26}+x_{1}x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{27}$
$\displaystyle\quad\quad\quad\ +x_{1}^{3}x_{2}x_{3}^{2}x_{4}x_{5}^{28}+x_{1}x_{2}^{3}x_{3}^{2}x_{4}x_{5}^{28}+x_{1}^{3}x_{2}x_{3}x_{4}^{2}x_{5}^{28}+x_{1}x_{2}^{3}x_{3}x_{4}^{2}x_{5}^{28}$
$\displaystyle\quad\quad\quad\ +x_{1}x_{2}x_{3}^{2}x_{4}^{3}x_{5}^{28}+x_{1}x_{2}x_{3}^{2}x_{4}x_{5}^{30}.$
Here $\psi$ is the homomorphism $\psi:(P_{5})_{15}\longrightarrow(P_{5})_{35},\,x_{1}^{e_{1}}\ldots x_{5}^{e_{5}}\longmapsto x_{1}^{2e_{1}+1}\ldots x_{5}^{2e_{5}+1},$ and the polynomial $q$ is determined as in Subsection 6.6 of [33, arXiv:1609.03006]. Re-verifying the above result by hand is also not too difficult. For the reader’s convenience, we also provide detailed output of our algorithm for the case $q=5,\,n=35$ at: https://doi.org/10.5281/zenodo.17620639.
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [34], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand using standard computational techniques from our collaborative work with Nguyen Sum (see [21, 22]). As a limitation of this approach, the result in [34] remains unverified by modern computer algebra systems such as OSCAR, Magma, or SageMath. In fact, verifying all the manual computations (now considered outdated) in [34] on a computer algebra system is very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [28]). This implies that the manual computational methods in [34], while not novel, become impractical at higher ranks, which explains why Singer’s Conjecture 1.1 remained open until our present work. As discussed above, we address this limitation for the $(q,n)=(6,36)$ case in Theorem 1.4, where we provide full computational verification via computer algebra systems.
Note 1.6.
Taking a different approach to Conjecture 1.1, Nguyen Huu Viet Hung [10] proposed the concept of a critical element within ${\rm Ext}^{q,*}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$. Specifically, a non-zero element $\mathbf{u}$ in ${\rm Ext}^{q,q+n}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is called critical, if it satisfies two conditions: (i) $\mu(2n+q)=q$, and (ii) the image of $\mathbf{u}$ under the classical squaring operation $Sq^{0}$ is zero.
It is well-established that $Sq^{0}$ is a monomorphism in positive stems of ${\rm Ext}_{\mathscr{A}}^{q,q+n}(\mathbb{F}_{2},\mathbb{F}_{2})$ for $q<5,$ thereby implying the absence of any critical element for $q<5.$ Remarkably, Hưng’s work [10, Theorem 5.9] states that Singer’s Conjecture 1.1 is not valid, if the algebraic transfer detects critical elements.
In [26], we proved that the non-zero element $D_{2}\in{\rm Ext}^{6,6+26}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is critical, but it is not in the image of $Tr_{6}(\mathbb{F}_{2}).$ Thus, the condition under which Hung’s work [10] would imply a negation of the conjecture was not met, and as we showed in [26], Conjecture 1.1 remains valid for bidegree $(6,6+26).$ This result, which was previously calculated entirely by hand, has been re-verified using the novel algorithm in the present work, yielding consistent results.
Additionally, in the case where the $\mathscr{A}$-module $\mathbb{F}_{2}\equiv\widetilde{H}^{*}\mathbb{S}^{0}$ is replaced by $\widetilde{H}^{*}\mathbb{R}P^{\infty}$, we have the non-zero element $\widehat{D_{2}}\in{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2}),$ and the Singer transfer is of the form
$Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}):(\mathbb{F}_{2}\otimes_{GL(q)}\mathcal{P}_{\mathscr{A}}(H_{*}\mathcal{V}^{q}\otimes\widetilde{H}_{*}\mathbb{R}P^{\infty}))_{n}\longrightarrow{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2}).$
Following Hưng [12, Theorem 2.1], if a critical element $\widehat{\mathbf{u}}\in{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ is in the image of the transfer $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}),$ then $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ is not a monomorphism. By [12, Theorem 2.2], the existence of a positive stem critical element $\widehat{\mathbf{u}}\in{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ in the image of the transfer $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ is equivalent to the existence of a positivestem critical element $\mathbf{u}$ in the image of the transfer $Tr_{q+1}(\mathbb{F}_{2}).$ If the existences happen, then both $Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$ and $Tr_{q+1}(\mathbb{F}_{2})$ are not injective.
We know that the algebraic Kahn-Priddy homomorphism $t_{*}$ defined by
$t_{*}:{\rm Ext}^{q,n+q}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})\longrightarrow{\rm Ext}^{q+1,n+q+1}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$
is a surjection in positive stems (see [12]). In particular, the restriction of $t_{*}$ maps ${\rm Im}(Tr_{q}(\widetilde{H}^{*}\mathbb{R}P^{\infty}))$ onto ${\rm Im}(Tr_{q+1}(\mathbb{F}_{2}))$. Hence, for $q=5$ and $n=26$, there exists a non-zero element $\widehat{D_{2}}$ in the image of the transfer
$Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty}):(\mathbb{F}_{2}\otimes_{GL(5)}\mathcal{P}_{\mathscr{A}}(H_{*}\mathcal{V}^{5}\otimes\widetilde{H}_{*}\mathbb{R}P^{\infty}))_{26}\longrightarrow{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$
with $\widehat{D_{2}}\in{\rm Ext}^{5,5+26}_{\mathscr{A}}(\widetilde{H}^{*}\mathbb{R}P^{\infty},\mathbb{F}_{2})$ and $t_{*}(\widehat{D_{2}})=D_{2}\in{\rm Ext}^{6,6+26}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$. Consequently, by a completely analogous argument to that used for the elements $\widehat{Ph_{1}}$ and $\widehat{Ph_{2}}$ in [12], it may be concluded that $\widehat{D_{2}}$ is also a critical element and is not in the image of $Tr_{5}(\widetilde{H}^{*}\mathbb{R}P^{\infty})$.
Theoretically, the approach to Singer’s Conjecture 1.1 via critical elements is promising. In practice, however, finding and characterizing these critical elements is computationally difficult.
For $q=6,n=36$, we see that the non-zero element $t\in\mathrm{Ext}^{6,6+36}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is not critical, since $\mu(2\cdot 36+6)=2<6$. However, we do not know whether this $t$ is in the image of $Tr_{6}(\mathbb{F}_{2})$ or not. Due to Theorem 1.4, we can propose the following.
Conjecture 1.7.
The non-zero element $t\in\mathrm{Ext}^{6,6+36}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})$ is detected by the sixth algebraic transfer $Tr_{6}(\mathbb{F}_{2})$.
It is known, by Chen [6], that the following element $\widetilde{t}$ is a representative of $t$:
$\widetilde{t}$ $=\lambda_{5}\bigl(\lambda_{9}\lambda_{3}\lambda_{5}\lambda_{7}^{2}+\lambda_{6}\lambda_{0}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{3}\lambda_{5}\lambda_{1}\lambda_{15}\lambda_{7}\bigr)$
$+\lambda_{3}\bigl(\lambda_{8}\lambda_{0}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{6}\lambda_{2}\lambda_{3}\lambda_{15}\lambda_{7}+\lambda_{5}\lambda_{9}\lambda_{5}\lambda_{7}^{2}+\lambda_{3}\lambda_{5}\lambda_{7}\lambda_{3}\lambda_{15}\bigr).$
Using this result together with Theorem 1.4 and our algorithm given in [29] for determining preimages in the lambda algebra, we hope that there will be an answer to Conjecture 1.7.
2 A few preliminaries
For substantiating our main result, namely Theorem 1.4, we recall underlying definitions and necessary ancillary homomorphisms. Extra specifics concerning these are obtainable through the works by [13, 29, 32].
As discussed in Section 1, our focus is on understanding both the behavior of the Singer algebraic transfer and Conjecture 1.1. In particular, the Singer conjecture is essential for studying the structure of the cohomology groups of the Steenrod algebra. To address this conjecture for the $q=6$ case, we need to explicitly determine the domain and codomain of the transfer map $Tr_{6}(\mathbb{F}_{2}).$ Remarkably, the domain of $Tr_{6}(\mathbb{F}_{2})$ is closely related to the problem of explicitly determining the dimension of the space $QP_{6}$ in positive degree $n.$ This issue is essentially about describing a minimal set of generators for the $\mathscr{A}$-module $P_{6}$, which is commonly referred to as the Peterson hit problem [20]. (For more perspectives on this remarkably difficult hit problem, we refer readers, for example, to our latest works [26].) Furthermore, it is well-known that the domain of $Tr_{q}(\mathbb{F}_{2})$ is dual to the $GL(q)$-invariant $[(QP_{q})_{n}]^{GL(q)}$ for any positive degree $n.$ Therefore, determining $\mathscr{A}$-generators for $P_{q}$ at degree $n$ stands as a crucial undertaking. Building on this relationship, we need to consider the following concepts.
Definition 2.1.
Let $\alpha_{j}(n)$ denote the $j$-th coefficient in the dyadic expansion of a positive integer $n$. This implies that $n$ can be written as $n=\sum_{j\geq 0}\alpha_{j}(n)2^{j}$, and each $\alpha_{j}(n)$ takes on values of 0 or 1. Consider a monomial $x=x_{1}^{a_{1}}x_{2}^{a_{2}}\ldots x_{q}^{a_{q}}\in P_{q}$. We define two associated sequences for $x$: $\omega(x)=(\omega_{1}(x),\omega_{2}(x),\ldots,\omega_{j}(x),\ldots),$ and $\sigma(x)=(a_{1},a_{2},\ldots,a_{q}),$ where $\omega_{j}(x)=\sum_{1\leq i\leq q}\alpha_{j-1}(a_{i})$ for $j\geq 1.$ Seeing that $\omega_{j}(x)\leq q$ for all $j.$ The sequences $\omega(x)$ (resp. $\sigma(x)$) are called the weight vector (resp. exponent vector) of $x.$
Vectors are compared using left lexicographic ordering.
We also want to emphasize that we can commence indexing for the weight vector $\omega(x)$ at zero, defining $\omega(x)=(\omega_{0}(x),\omega_{1}(x),\omega_{2}(x),\ldots,\omega_{j}(x),\ldots),$ where $\omega_{j}(x)=\sum_{1\leq i\leq q}\alpha_{j}(a_{i}),\,j\geq 0.$ However, in our view, following Definition 2.3 below concerning the comparison between two monomials related to weight vector and exponent vector, we believe it is advantageous to index the weight vector $\omega(x)$ starting at 1, akin to indexing the exponent vector, to facilitate comparison between two monomials.
For a weight vector $\omega=(\omega_{1},\omega_{2},\ldots,\omega_{j},0,0,\ldots,0),$ we define $\deg\omega=\sum\limits_{j\geq 1}2^{j-1}\omega_{j}.$ Denote by $P_{q}(\omega)$ the subspace of $P_{q}$ spanned by all monomials $x\in P_{q}$ such that $\deg x=\deg\omega,\ \omega(x)\leq\omega,$ and by $P_{q}^{-}(\omega)$ the subspace of $P_{q}(\omega)$ spanned by all monomials $x$ such that $\omega(x)<\omega.$
Definition 2.2.
Assume both $f$ and $g$ are homogeneous polynomials in $P_{q}$ such that $\deg(f)=\deg(g).$ The following binary relation $\equiv_{\omega}$” can be readily identified as an equivalence relation on $P_{q}$:
$f\equiv_{\omega}g\ \mbox{if and only if}\ (f+g)\in\mathscr{A}^{>0}\cdot P_{q}+P_{q}^{-}(\omega).$
If we denote $QP_{q}(\omega)$ as the quotient of the equivalence relation $\equiv_{\omega},$ then
$QP_{q}(\omega)=P_{5}(\omega)/((\mathscr{A}^{>0}\cdot P_{q}\cap P_{q}(\omega))+P_{q}^{-}(\omega)).$
Furthermore, as is well known [36, 37], $QP_{q}(\omega)$ is also a $GL(q)$-module.
From now on, if $f$ is a polynomial in $f\in P_{q}(\omega),$ then we denote by $[f]_{\omega}$ the class in $QP_{q}(\omega)$ represented by $f.$ For a set $S\subset P_{q}(\omega),$ denote by $[S]_{\omega}=\{[f]_{\omega}\in QP_{q}(\omega):\,f\in S\}\subset QP_{q}(\omega).$
Definition 2.3.
Given monomials $x$ and $y$ in $P_{q}$ with the same degree, the relation $y<x$ is defined by the condition that either $\omega(y)<\omega(x)$ or $\omega(x)=\omega(y)$ and $\sigma(y)<\sigma(x).$
Definition 2.4.
  1. (i)
    A monomial $x\in P_{q}$ is said to be inadmissible if there exist monomials $y_{1},y_{2},\ldots,y_{m}$ such that $\deg(x)=\deg(y_{j})$ and $y_{j}<x$ for $1\leq j\leq m$ and
    $x+\sum\limits_{1\leq j\leq m}y_{j}\in\mathscr{A}^{>0}\cdot P_{q}.$
  2. (ii)
    A monomial $x\in P_{5}$ is said to be admissible if it is not inadmissible.
Thus, it can be observed that $(QP_{q})_{n}$ is a $\mathbb{F}_{2}$-vector space, with its basis being composed of the classes represented by the admissible monomials in $(P_{q})_{n}$. From now on, we denote by $\mathbf{Ad}_{q}(\omega)$ the collection of all admissible monomials of degree $n$ in $P_{q}(\omega)$.
According to [36], we have an isomorphism
$(QP_{q})_{n}\cong\bigoplus\limits_{{\rm deg}\,\omega=n}QP_{q}(\omega).$
We refer the reader to our work [29] for a detailed proof of this result.
Definition 2.5.
For $1\leq j\leq q,$ we define the $\mathscr{A}$-homomorphism $\rho_{j}:P_{q}\to P_{q}$ by its action on the variables $\{x_{1},\dots,x_{q}\}$. The definition is split into two cases.
  • 1.
    Adjacent transpositions ($1\leq j\leq q-1$): The operator $\rho_{j}$ swaps the adjacent variables $x_{j}$ and $x_{j+1}$ and fixes all others:
    $\rho_{j}(x_{i})=\begin{cases}x_{j+1}&\text{if }i=j\\ x_{j}&\text{if }i=j+1\\ x_{i}&\text{otherwise}.\end{cases}$
  • 2.
    A transvection ($j=q$): The operator $\rho_{q}$ adds the variable $x_{q-1}$ to $x_{q}$ and fixes all others:
    $\rho_{q}(x_{i})=\begin{cases}x_{q}+x_{q-1}&\text{if }i=q\\ x_{i}&\text{if }i<q.\end{cases}$
The action of any $\rho_{j}$ is extended to all polynomials in $P_{q}$ by the property that it is an algebra homomorphism. Since every permutation is a product of transpositions, and every transposition is a product of adjacent transpositions (the operators $\rho_{j}$ for $j<q$), the set $\{\rho_{1},\dots,\rho_{q-1}\}$ generates the entire symmetric group $\Sigma_{q}\subset GL(q).$ Then, the general linear group $GL(q)$ is generated by the set of operators $\{\rho_{j}\mid 1\leq j\leq q\}.$
Let $[u]_{\omega}$ be a class in $QP_{q}(\omega)$ represented by a homogeneous polynomial $u\in P_{q}(\omega)$.
  • 1.
    The class $[u]_{\omega}$ is $\Sigma_{q}$-invariant if and only if it is invariant under the action of all adjacent transpositions:
    $\rho_{j}(u)+u\equiv_{\omega}0\quad\text{for all }j\in\{1,\dots,q-1\}.$
  • 2.
    The class $[u]_{\omega}$ is $GL(q)$-invariant if and only if it is $\Sigma_{q}$-invariant and is also invariant under the action of the transvection $\rho_{q}.$ This is equivalent to the single, comprehensive condition:
    $\rho_{j}(u)+u\equiv_{\omega}0\quad\text{for all }j\in\{1,\dots,q\}.$
3 Proof of Theorem 1.4
It is worth noting that the proof of Theorem 1.4 is presented via our algorithms in SageMath and OSCAR to obviate the need for transcribing unnecessary manual computations. These computations are already detailed in the algorithm’s output (see Note 3.5(B)). Consequently, rather than detailing such lengthy calculations, we sketch the proof based on the construction of our new algorithm.
Remark 3.1.
In order to prove Theorem 1.4, we use the Kameko homomorphism [13]:
$\displaystyle(\widetilde{Sq}^{0}_{*})_{(q,2n+q)}:(QP_{q})_{2n+q}$ $\displaystyle\longrightarrow(QP_{q})_{n},$
$\displaystyle\mbox{[}x_{1}^{a_{1}}x_{2}^{a_{2}}\ldots x_{q}^{a_{q}}\mbox{]}$ $\displaystyle\longmapsto\left\{\begin{array}[]{ll}\mbox{[}x_{1}^{\frac{a_{1}-1}{2}}x_{2}^{\frac{a_{2}-1}{2}}\ldots x_{q}^{\frac{a_{q}-1}{2}}\mbox{]}&\text{if $\displaystyle a_{1},a_{2},\ldots,a_{q}$ odd},\\ 0&\text{otherwise}.\end{array}\right$.
It is well-known that $(\widetilde{Sq}^{0}_{*})_{(q,2n+q)}$ is surjective. Hence,
$\dim(QP_{q})_{2n+q}=\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,2n+q)})+\dim(QP_{q})_{n}.$
With $q=6$ and $n=36,$ we have
$\dim(QP_{6})_{36}=\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})+\dim(QP_{6})_{15}.$
In [Phuc4], we showed that $\dim(QP_{6})_{15}=2184.$ So, we need only to determine $\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ to deduce the dimensional result for $(QP_{6})_{36}.$ And from that, based on a basis for $(QP_{6})_{36}$ and the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 6,$ we can explicitly compute the dimension and basis for the invariant space $[(QP_{6})_{36}]^{GL(6)}.$
We notice that computing by hand an explicit basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ is a hard and error-prone task, due to the growing number of monomials as the number of variables and degrees increases. Manual computation can typically be controlled in cases where the degree is not too large and the increase in the number of monomials is manageable (for instance, one can see some works by the author [25, 26], the author and Nguyen Sum [21], and Nguyen Sum [32] to understand the specific manual computation methods). Therefore, in recent works [27, 28, 29], we have developed computational programs to explicitly compute the dimension of $(QP_{q})_{n}$ and the invariant $[(QP_{q})_{n}]^{GL(q)}.$ These algorithmic programs allow us to display detailed computations as readers have become familiar with in our previous works, and some other authors.
Now, with $q=6$ and $n=15,$ by using our algorithm in [29], we obtain the following:
  • (i)
    $(QP_{6})_{15}\cong\bigoplus_{1\leq i\leq 7}QP_{6}(\omega_{(i)}),$ where
    $\displaystyle\omega_{(1)}:=(1,1,1,1),\ \ \omega_{(2)}:=(1,1,3),\ \ \omega_{(3)}:=(1,3,2),\ \ \omega_{(4)}:=(3,2,2),$
    $\displaystyle\omega_{(5)}:=(3,4,1),\ \ \omega_{(6)}:=(5,3,1),\ \ \omega_{(7)}:=(5,5).$
  • (ii)
    We have
    $\dim QP_{6}(\omega_{(i)})=\left\{\begin{array}[]{ll}56&\mbox{if $\displaystyle i=1$},\\ 6&\mbox{if $\displaystyle i=2$},\\ 1&\mbox{if $\displaystyle i=3$},\\ 1176&\mbox{if $\displaystyle i=4$},\\ 384&\mbox{if $\displaystyle i=5$},\\ 540&\mbox{if $\displaystyle i=6$},\\ 21&\mbox{if $\displaystyle i=7$}.\end{array}\right$.
  • (iii)
    We have
    $\dim[QP_{6}(\omega_{(i)})]^{GL(6)}=\left\{\begin{array}[]{ll}1&\mbox{if $\displaystyle i=3$},\\ 0&\mbox{otherwise}.\end{array}\right$.
    Furthermore, $[QP_{6}(\omega_{(3)})]^{GL(6)}=\mathbb{F}_{2}\cdot[x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}]_{\omega_{(3)}}.$
Assume that $g\in P_{6}$ such that $[g]\in[(QP_{6})_{15}]^{GL(6)}.$ Then,
$g\equiv\beta\cdot x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+\sum_{x\in\mathbf{Ad}_{6}(\omega_{(1)})\cup\mathbf{Ad}_{6}(\omega_{(2)}))}\beta_{x}\cdot x,\ \ \beta,\,\beta_{x}\in\mathbb{F}_{2},$
where $\mathbf{Ad}_{6}(\omega_{(1)})$ is the set consisting of the following 56 admissible monomials:
$\displaystyle[1].\,x_{3}x_{4}^{2}x_{5}^{4}x_{6}^{8},$ $\displaystyle\quad[2].\,x_{2}x_{4}^{2}x_{5}^{4}x_{6}^{8},$ $\displaystyle\quad[3].\,x_{2}x_{3}^{2}x_{5}^{4}x_{6}^{8},$
$\displaystyle[4].\,x_{2}x_{3}^{2}x_{4}^{4}x_{6}^{8},$ $\displaystyle\quad[5].\,x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{8},$ $\displaystyle\quad[6].\,x_{1}x_{4}^{2}x_{5}^{4}x_{6}^{8},$
$\displaystyle[7].\,x_{1}x_{3}^{2}x_{5}^{4}x_{6}^{8},$ $\displaystyle\quad[8].\,x_{1}x_{3}^{2}x_{4}^{4}x_{6}^{8},$ $\displaystyle\quad[9].\,x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{8},$
$\displaystyle[10].\,x_{1}x_{2}^{2}x_{5}^{4}x_{6}^{8},$ $\displaystyle\quad[11].\,x_{1}x_{2}^{2}x_{4}^{4}x_{6}^{8},$ $\displaystyle\quad[12].\,x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{8},$
$\displaystyle[13].\,x_{1}x_{2}^{2}x_{3}^{4}x_{6}^{8},$ $\displaystyle\quad[14].\,x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{8},$ $\displaystyle\quad[15].\,x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{8},$
$\displaystyle[16].\,x_{4}x_{5}^{2}x_{6}^{12},$ $\displaystyle\quad[17].\,x_{3}x_{5}^{2}x_{6}^{12},$ $\displaystyle\quad[18].\,x_{3}x_{4}^{2}x_{6}^{12},$
$\displaystyle[19].\,x_{3}x_{4}^{2}x_{5}^{12},$ $\displaystyle\quad[20].\,x_{2}x_{5}^{2}x_{6}^{12},$ $\displaystyle\quad[21].\,x_{2}x_{4}^{2}x_{6}^{12},$
$\displaystyle[22].\,x_{2}x_{4}^{2}x_{5}^{12},$ $\displaystyle\quad[23].\,x_{2}x_{3}^{2}x_{6}^{12},$ $\displaystyle\quad[24].\,x_{2}x_{3}^{2}x_{5}^{12},$
$\displaystyle[25].\,x_{2}x_{3}^{2}x_{4}^{12},$ $\displaystyle\quad[26].\,x_{1}x_{5}^{2}x_{6}^{12},$ $\displaystyle\quad[27].\,x_{1}x_{4}^{2}x_{6}^{12},$
$\displaystyle[28].\,x_{1}x_{4}^{2}x_{5}^{12},$ $\displaystyle\quad[29].\,x_{1}x_{3}^{2}x_{6}^{12},$ $\displaystyle\quad[30].\,x_{1}x_{3}^{2}x_{5}^{12},$
$\displaystyle[31].\,x_{1}x_{3}^{2}x_{4}^{12},$ $\displaystyle\quad[32].\,x_{1}x_{2}^{2}x_{6}^{12},$ $\displaystyle\quad[33].\,x_{1}x_{2}^{2}x_{5}^{12},$
$\displaystyle[34].\,x_{1}x_{2}^{2}x_{4}^{12},$ $\displaystyle\quad[35].\,x_{1}x_{2}^{2}x_{3}^{12},$ $\displaystyle\quad[36].\,x_{5}x_{6}^{14},$
$\displaystyle[37].\,x_{4}x_{6}^{14},$ $\displaystyle\quad[38].\,x_{4}x_{5}^{14},$ $\displaystyle\quad[39].\,x_{3}x_{6}^{14},$
$\displaystyle[40].\,x_{3}x_{5}^{14},$ $\displaystyle\quad[41].\,x_{3}x_{4}^{14},$ $\displaystyle\quad[42].\,x_{2}x_{6}^{14},$
$\displaystyle[43].\,x_{2}x_{5}^{14},$ $\displaystyle\quad[44].\,x_{2}x_{4}^{14},$ $\displaystyle\quad[45].\,x_{2}x_{3}^{14},$
$\displaystyle[46].\,x_{1}x_{6}^{14},$ $\displaystyle\quad[47].\,x_{1}x_{5}^{14},$ $\displaystyle\quad[48].\,x_{1}x_{4}^{14},$
$\displaystyle[49].\,x_{1}x_{3}^{14},$ $\displaystyle\quad[50].\,x_{1}x_{2}^{14},$ $\displaystyle\quad[51].\,x_{6}^{15},$
$\displaystyle[52].\,x_{5}^{15},$ $\displaystyle\quad[53].\,x_{4}^{15},$ $\displaystyle\quad[54].\,x_{3}^{15},$
$\displaystyle[55].\,x_{2}^{15},$ $\displaystyle\quad[56].\,x_{1}^{15},$
and $\mathbf{Ad}_{6}(\omega_{(2)})$ is the set consisting of the following 6 admissible monomials:
$\displaystyle[57].\,x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4},$ $\displaystyle\quad[58].\,x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4},$ $\displaystyle\quad[59].\,x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4},$
$\displaystyle[60].\,x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{4}x_{6}^{4},$ $\displaystyle\quad[61].\,x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{6}^{4},$ $\displaystyle\quad[62].\,x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{5}^{4}.$
By a direct computation using the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 5,$ we find that $\rho_{j}(g)\equiv g$ if and only if
$g\equiv\beta\cdot x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+\sum_{1\leq i\leq 5}\beta_{i}h_{i},$
where
$\displaystyle h_{1}$ $\displaystyle=x_{1}x_{2}^{14}+x_{1}x_{3}^{14}+x_{2}x_{3}^{14}+x_{1}x_{4}^{14}$
$\displaystyle\quad+x_{2}x_{4}^{14}+x_{3}x_{4}^{14}+x_{1}x_{5}^{14}+x_{2}x_{5}^{14}$
$\displaystyle\quad+x_{3}x_{5}^{14}+x_{4}x_{5}^{14}+x_{1}x_{6}^{14}+x_{2}x_{6}^{14}$
$\displaystyle\quad+x_{3}x_{6}^{14}+x_{4}x_{6}^{14}+x_{5}x_{6}^{14},$
$\displaystyle h_{2}$ $\displaystyle=x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{8}+x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{8}+x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{8}+x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{8}$
$\displaystyle\quad+x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{8}+x_{1}x_{2}^{2}x_{3}^{4}x_{6}^{8}+x_{1}x_{2}^{2}x_{4}^{4}x_{6}^{8}+x_{1}x_{3}^{2}x_{4}^{4}x_{6}^{8}$
$\displaystyle\quad+x_{2}x_{3}^{2}x_{4}^{4}x_{6}^{8}+x_{1}x_{2}^{2}x_{5}^{4}x_{6}^{8}+x_{1}x_{3}^{2}x_{5}^{4}x_{6}^{8}+x_{2}x_{3}^{2}x_{5}^{4}x_{6}^{8}$
$\displaystyle\quad+x_{1}x_{4}^{2}x_{5}^{4}x_{6}^{8}+x_{2}x_{4}^{2}x_{5}^{4}x_{6}^{8}+x_{3}x_{4}^{2}x_{5}^{4}x_{6}^{8},$
$\displaystyle h_{3}$ $\displaystyle=x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{5}^{4}+x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{6}^{4}+x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{4}x_{6}^{4}+x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4}$
$\displaystyle\quad+x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4}+x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4},$
$\displaystyle h_{4}$ $\displaystyle=x_{1}^{15}+x_{2}^{15}+x_{3}^{15}+x_{4}^{15}+x_{5}^{15}+x_{6}^{15},$
$\displaystyle h_{5}$ $\displaystyle=x_{1}x_{2}^{2}x_{3}^{12}+x_{1}x_{2}^{2}x_{4}^{12}+x_{1}x_{3}^{2}x_{4}^{12}+x_{2}x_{3}^{2}x_{4}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{2}x_{5}^{12}+x_{1}x_{3}^{2}x_{5}^{12}+x_{2}x_{3}^{2}x_{5}^{12}+x_{1}x_{4}^{2}x_{5}^{12}$
$\displaystyle\quad+x_{2}x_{4}^{2}x_{5}^{12}+x_{3}x_{4}^{2}x_{5}^{12}+x_{1}x_{2}^{2}x_{6}^{12}+x_{1}x_{3}^{2}x_{6}^{12}$
$\displaystyle\quad+x_{2}x_{3}^{2}x_{6}^{12}+x_{1}x_{4}^{2}x_{6}^{12}+x_{2}x_{4}^{2}x_{6}^{12}+x_{3}x_{4}^{2}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{5}^{2}x_{6}^{12}+x_{2}x_{5}^{2}x_{6}^{12}+x_{3}x_{5}^{2}x_{6}^{12}+x_{4}x_{5}^{2}x_{6}^{12}.$
Finally, based on the relation $\rho_{6}(g)\equiv g,$ we obtain $\beta=\beta_{i}$ for all $i,\,1\leq i\leq 5.$ Consequently,
$g\equiv\beta\bigg(x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+\sum_{1\leq j\leq 5}h_{j}\bigg),\,\beta\in\mathbb{F}_{2}.$
Thus, we obtain the following:
Proposition 3.2.
We have
$\dim[(QP_{6})_{15}]^{GL(6)}=1,\ \ \mbox{and}\ \ [(QP_{6})_{15}]^{GL(6)}=\mathbb{F}_{2}\cdot[\xi],$
where
$\displaystyle\xi$ $\displaystyle=x_{1}^{15}+x_{1}x_{2}^{14}+x_{2}^{15}+x_{1}x_{2}^{2}x_{3}^{12}$
$\displaystyle\quad+x_{1}x_{3}^{14}+x_{2}x_{3}^{14}+x_{3}^{15}+x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{8}$
$\displaystyle\quad+x_{1}x_{2}^{2}x_{4}^{12}+x_{1}x_{3}^{2}x_{4}^{12}+x_{2}x_{3}^{2}x_{4}^{12}+x_{1}x_{4}^{14}$
$\displaystyle\quad+x_{2}x_{4}^{14}+x_{3}x_{4}^{14}+x_{4}^{15}+x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{5}^{4}$
$\displaystyle\quad+x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{8}+x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{8}+x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{8}+x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{8}$
$\displaystyle\quad+x_{1}x_{2}^{2}x_{5}^{12}+x_{1}x_{3}^{2}x_{5}^{12}+x_{2}x_{3}^{2}x_{5}^{12}+x_{1}x_{4}^{2}x_{5}^{12}$
$\displaystyle\quad+x_{2}x_{4}^{2}x_{5}^{12}+x_{3}x_{4}^{2}x_{5}^{12}+x_{1}x_{5}^{14}+x_{2}x_{5}^{14}$
$\displaystyle\quad+x_{3}x_{5}^{14}+x_{4}x_{5}^{14}+x_{5}^{15}+x_{1}x_{2}^{2}x_{3}^{4}x_{4}^{4}x_{6}^{4}$
$\displaystyle\quad+x_{1}x_{2}^{2}x_{3}^{4}x_{5}^{4}x_{6}^{4}+x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+x_{1}x_{2}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4}+x_{1}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4}$
$\displaystyle\quad+x_{2}x_{3}^{2}x_{4}^{4}x_{5}^{4}x_{6}^{4}+x_{1}x_{2}^{2}x_{3}^{4}x_{6}^{8}+x_{1}x_{2}^{2}x_{4}^{4}x_{6}^{8}+x_{1}x_{3}^{2}x_{4}^{4}x_{6}^{8}$
$\displaystyle\quad+x_{2}x_{3}^{2}x_{4}^{4}x_{6}^{8}+x_{1}x_{2}^{2}x_{5}^{4}x_{6}^{8}+x_{1}x_{3}^{2}x_{5}^{4}x_{6}^{8}+x_{2}x_{3}^{2}x_{5}^{4}x_{6}^{8}$
$\displaystyle\quad+x_{1}x_{4}^{2}x_{5}^{4}x_{6}^{8}+x_{2}x_{4}^{2}x_{5}^{4}x_{6}^{8}+x_{3}x_{4}^{2}x_{5}^{4}x_{6}^{8}+x_{1}x_{2}^{2}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{3}^{2}x_{6}^{12}+x_{2}x_{3}^{2}x_{6}^{12}+x_{1}x_{4}^{2}x_{6}^{12}+x_{2}x_{4}^{2}x_{6}^{12}$
$\displaystyle\quad+x_{3}x_{4}^{2}x_{6}^{12}+x_{1}x_{5}^{2}x_{6}^{12}+x_{2}x_{5}^{2}x_{6}^{12}+x_{3}x_{5}^{2}x_{6}^{12}$
$\displaystyle\quad+x_{4}x_{5}^{2}x_{6}^{12}+x_{1}x_{6}^{14}+x_{2}x_{6}^{14}+x_{3}x_{6}^{14}$
$\displaystyle\quad+x_{4}x_{6}^{14}+x_{5}x_{6}^{14}+x_{6}^{15}.$
Detailed computations for this result are shown in the output of our algorithm in Note 3.5(A) below.
In the next step, we will explicitly compute the basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ and the $GL(6)$-invariant $[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}.$ As mentioned above, computing these spaces by hand seems infeasible and error-prone due to the prohibitively large number of input monomials (specifically, by the formula in [28], $\dim(P_{6})_{36}=\binom{36+(6-1)}{6-1}=749,398$). To overcome this difficulty, we will construct an algorithmic program implemented in the computer algebra system OSCAR [39] that allows us to explicitly determine the basis of the spaces ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$ and of their $GL(q)$-invariant subspaces for any $q$ and $n$ satisfying $n-q$ even. Based on the previously obtained results for the $(GL(q)$-invariants of the target space of the Kameko homomorphism $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ (computed via our algorithm in [29]), we proceed as follows: We first compute the $(GL(q)$-invariants in degree $(n-q)/2$ for the target of $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ by the method of [29]. These target invariants are then used as seeds: we apply the inverse Kameko lift $\psi:\ (P_{q})_{\frac{n-q}{2}}\longrightarrow(P_{q})_{n},\ \ x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}\longmapsto x_{1}^{2e_{1}+1}\cdots x_{q}^{2e_{q}+1},$ and, on the subset of admissible coordinates contained in ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$, we solve the stacked linear systems enforcing $(\rho_{j}-\mathrm{Id})\,f\equiv 0\quad(j=1,\dots,q-1)$ and $(\rho_{q}-\mathrm{Id})\,f\equiv 0,$ thereby correcting the lifts to genuine $GL(q)$-invariants. In parallel, within the kernel itself we perform a weightwise computation of $\Sigma_{q}$- and $GL(q)$-invariants and then apply a largest-weight correction to non-zero $GL(q)$-invariants. Below we construct in detail our algorithm as sketched.
$\displaystyle\trianglerightIntegers$ $q\geq 1$, $n\geq 0$ with $n\equiv q\pmod{2}$; base field $\mathbb{F}_{2}$.
$\displaystyle\trianglerightA$ basis of $GL(q)$-invariants inside $(QP_{q})_{n}$ obtained by: streaming hit elimination, Kameko kernel, weightwise $\Sigma_{q}$$GL(q).$
[…]
$\displaystyle\triangleright$  $m\leftarrow\max_{i}a_{i}$; if $m=0$ return empty vector
$\displaystyle\triangleright$  $t\leftarrow m$; $L\leftarrow 0$;
$\displaystyle\triangleright$  while $t>0$ do $t\leftarrow\lfloor t/2\rfloor$; $L\leftarrow L+1$
$\displaystyle\triangleright$  end while
$\displaystyle\triangleright$  for $b=0,\dots,L-1$ do $\omega_{b+1}\leftarrow\sum_{i=1}^{q}\big((a_{i}\div 2^{b})\bmod 2\big)$
$\displaystyle\triangleright$  end for
$\displaystyle\triangleright$  return $\omega=(\omega_{1},\dots,\omega_{L})$
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ KamekoImageExps($\mathbf{a}$)
$\displaystyle\triangleright$  if some $a_{i}$ is even then return None
$\displaystyle\triangleright$  else return $\big((a_{1}-1)/2,\dots,(a_{q}-1)/2\big)$
$\displaystyle\triangleright$  end if
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ SqOnMono($k,\;x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}$) $\triangleright$ Cartan + Lucas mod 2
$\displaystyle\triangleright$  if $k=0$ then return $x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}$
$\displaystyle\triangleright$  end if
$\displaystyle\triangleright$  Pick first $j$ with $e_{j}>0$; write $x_{j}^{e_{j}}\cdot M^{\prime}$; $Sq^{k}(M)=\sum_{i=0}^{k}\binom{e_{j}}{i}x_{j}^{e_{j}+i}\,Sq^{k-i}(M^{\prime})$ over $\mathbb{F}_{2}$
$\displaystyle\triangleright$  return result (with memoization)
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ HitColumnExps($\mathbf{a},\;k$)
$\displaystyle\triangleright$  $M\leftarrow x_{1}^{a_{1}}\cdots x_{q}^{a_{q}}$; $S\leftarrow Sq^{k}(M)$
$\displaystyle\triangleright$  Collect exponent tuples of monomials in $S$ with odd parity (mod 2), sorted
$\displaystyle\triangleright$  return list of exponent tuples
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ ExpsEnum($q,n$) $\triangleright$ All $\mathbf{a}\in\mathbb{N}^{q}$ with $\sum a_{i}=n$
$\displaystyle\triangleright$  return the standard stars-and-bars enumeration
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ BuildDegSpaceOnline($q,n$) $\triangleright$ Stream + ONLINE elimination in degree $n$
$\displaystyle\triangleright$  $\mathcal{E}\leftarrow$ ExpsEnum($q,n$); sort $\mathcal{E}$ by $(\omega(\mathbf{a}),\mathbf{a})$ lexicographic
$\displaystyle\triangleright$  Make dictionary $\texttt{idx}:\mathcal{E}\to\{1,\dots,|\mathcal{E}|\}$; $\texttt{pivotmap}\leftarrow\emptyset$
$\displaystyle\triangleright$  for $p=0$ while $2^{p}\leq n$ do
$\displaystyle\triangleright$   $k\_{{\rm op}}\leftarrow 2^{p}$; $n_{g}\leftarrow n-k\_{{\rm op}}$
$\displaystyle\triangleright$   for all $\mathbf{b}\in$ ExpsEnum($q,n_{g}$) do $\triangleright$ stream
$\displaystyle\triangleright$     $R\leftarrow$ HitColumnExps($\mathbf{b},k\_{{\rm op}}$)
$\displaystyle\triangleright$     Map each $\mathbf{r}\in R$ to row index $r=\texttt{idx}(\mathbf{r})$ (drop if missing)
$\displaystyle\triangleright$     Reduce the sorted row-list online by XOR against pivotmap (keep new pivot if any)
$\displaystyle\triangleright$   end for
$\displaystyle\triangleright$  end for
$\displaystyle\triangleright$  $S\_\text{pivots}\leftarrow$ keys of pivotmap; admissible indices $A\leftarrow\{1,\dots,|\mathcal{E}|\}\setminus S\_\text{pivots}$
$\displaystyle\triangleright$  return $\mathsf{DS}(q,n)$ with fields: $\mathcal{E}$, idx, admissible exponents $\{\mathbf{a}_{i}\}_{i\in A}$, and the online reduction data
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ ReduceRowToAdmissible($r,\mathsf{DS}$) $\triangleright$ Global row $\to$ admissible positions
$\displaystyle\triangleright$  XOR-reduce $[r]$ by pivotmap until no pivot hits; map survivors to positions in admissible list
$\displaystyle\triangleright$  return sorted position-list
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ BuildKamekoBitMat($\mathsf{DS\_src},\mathsf{DS\_tgt}$)
$\displaystyle\triangleright$  Make bit-matrix $L$ of size $(\dim\text{adm tgt})\times(\dim\text{adm src})$
$\displaystyle\triangleright$  for each source admissible exponent $\mathbf{a}$ with column $c$ do
$\displaystyle\triangleright$   $\mathbf{u}\leftarrow$ KamekoImageExps($\mathbf{a}$);
$\displaystyle\triangleright$   if $\mathbf{u}=\textsc{None}$ then continue
$\displaystyle\triangleright$   end if
$\displaystyle\triangleright$   $r\leftarrow\texttt{idx}_{\rm tgt}(\mathbf{u})$ (skip if missing); rows $\leftarrow$ ReduceRowToAdmissible($r,\mathsf{DS\_tgt}$)
$\displaystyle\triangleright$   Set the bits $L[\text{rows},c]\leftarrow 1$
$\displaystyle\triangleright$  end for
$\displaystyle\triangleright$  return $L$
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ NullspaceGFTwo(bit-matrix $M$)
$\displaystyle\triangleright$  Perform bit-packed Gaussian elimination over $\mathbb{F}_{2}$
$\displaystyle\triangleright$  return (rank, list of nullspace basis vectors)
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ ApplyRho($j,\,x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}$)
$\displaystyle\triangleright$  if $1\leq j<q$ then swap $x_{j},x_{j+1}$;
$\displaystyle\triangleright$  else if $j=q$ then send $x_{q}\mapsto x_{q}+x_{q-1}$;
$\displaystyle\triangleright$  else return identity
$\displaystyle\triangleright$  end if
$\displaystyle\triangleright$  Extend multiplicatively to polynomials
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ DecomposeToEntries($f,\mathsf{DS}$)
$\displaystyle\triangleright$  Write $f$ as $\mathbb{F}_{2}$-sum of monomials; map each to global row, reduce to admissible positions (with parity)
$\displaystyle\triangleright$  return sorted list of admissible positions
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ PrecomputeRhoRows($\mathsf{DS}$) $\triangleright$ Rows of $(\rho_{j}-\mathrm{Id})$ on each admissible basis element
$\displaystyle\triangleright$  for $j=1,\dots,q$ do
$\displaystyle\triangleright$   for each admissible mono $u_{i}$ do
[…]
$\displaystyle\triangleright$   end for
$\displaystyle\triangleright$  end for
$\displaystyle\trianglerightend$ function
$\displaystyle\trianglerightfunction$ $\Sigma_{q}/GL(q)$-On-Kernel-Weight($\mathsf{DS},\;\ker L,\;\mathcal{I}_{w}$)
$\displaystyle\triangleright$  $\mathcal{I}_{w}$: indices of admissible monomials of fixed weight $\omega$ that appear in some kernel vector
$\displaystyle\triangleright$  Let $\{u_{1},\dots,u_{N_{w}}\}$ be those monomials; pick kernel columns that meet $\mathcal{I}_{w}$
$\displaystyle\triangleright$  ($\Sigma_{q}$-stage) Build stacked matrix of $(\rho_{j}-\mathrm{Id})\sum_{i}\gamma_{i}u_{i}$ for $j=1,\dots,q-1$; find nullspace
$\displaystyle\triangleright$  Obtain $\Sigma_{q}$-basis $\{\sum_{i}\gamma^{(t)}_{i}u_{i}\}_{t}$
$\displaystyle\triangleright$  ($GL(q)$-stage) Build matrix of $(\rho_{q}-\mathrm{Id})\sum_{t}\beta_{t}(\Sigma\text{-basis})_{t}$; find nullspace
$\displaystyle\triangleright$  Obtain weightwise $GL(q)$-invariants $\{\sum_{i}\lambda^{(s)}_{i}u_{i}\}_{s}$
$\displaystyle\triangleright$  Note (diagnostic only): grouping coordinates by $\gamma$$\beta$-signature” is for reporting structure of solutions and does not affect any nullspace computation.
$\displaystyle\triangleright$  return ($\Sigma_{q}$-basis, $GL(q)$-basis) in this weight
$\displaystyle\trianglerightend$ function
$\displaystyle\triangleright$
$\displaystyle\trianglerightprocedure$ RunAll($q,n$) $\triangleright$ Main orchestration
$\displaystyle\triangleright$  require $n\equiv q\pmod{2}$; $n_{\rm tgt}\leftarrow(n-q)/2$
$\displaystyle\triangleright$  [Step 1] $\mathsf{DS\_src}\leftarrow$ BuildDegSpaceOnline($q,d$) $\triangleright$ $QP_{q}$-basis by streaming hit elimination
$\displaystyle\triangleright$  [Step 2] $\mathsf{DS\_tgt}\leftarrow$ BuildDegSpaceOnline($q,n_{\rm tgt}$)
$\displaystyle\triangleright$  [Step 3] $L\leftarrow$ BuildKamekoBitMat($\mathsf{DS\_src},\mathsf{DS\_tgt}$); $(\mathrm{rk},\ker L)\leftarrow$ NullspaceGFTwo($L$)
$\displaystyle\triangleright$  PrecomputeRhoRows($\mathsf{DS\_src}$)
$\displaystyle\triangleright$  Extract kernel support indices $\mathcal{K}\subset$ admissible positions of source; group by weights $\omega$
$\displaystyle\triangleright$  for each weight $\omega$ having $\mathcal{I}_{\omega}:=\mathcal{K}\cap\{\text{weight}=\omega\}\neq\emptyset$ do
$\displaystyle\triangleright$   $(\Sigma_{q}[\omega],\,GL(q)[\omega])\leftarrow$ $\Sigma_{q}/GL(q)$-On-Kernel-Weight($\mathsf{DS\_src},\ker L,\mathcal{I}_{\omega}$)
$\displaystyle\triangleright$  end for
$\displaystyle\triangleright$  [Step 4] Correction inside kernel (largest weight with $GL(q)\neq 0$):
$\displaystyle\triangleright$  if all $GL(q)[\omega]$ are empty then report $GL(q)$-invariants in kernel $=0$
$\displaystyle\triangleright$  else
$\displaystyle\triangleright$   pick $\omega^{\star}=\max\{\omega:GL(q)[\omega]\neq\emptyset\}$; set $\mathcal{L}:=\{i\in\mathcal{K}:\text{weight}(i)<\omega^{\star}\}$
$\displaystyle\triangleright$   for each $g_{\max}\in GL(q)[\omega^{\star}]$ do
$\displaystyle\triangleright$     (Stage 1) Solve on subset $\mathcal{K}$ for
$\phi=\gamma\cdot g_{\max}+\sum_{t\in\mathcal{L}}\beta_{t}\,u_{t}\quad\text{with}\quad(\rho_{j}-\mathrm{Id})\phi\equiv 0,\;j=1,\dots,q-1,$
$\displaystyle\triangleright$     i.e. build stacked matrix on $\mathcal{K}$ and take nullspace to get a basis $\{\phi_{s}\}_{s}$
$\displaystyle\triangleright$     (Stage 2) Solve $\sum_{s}\lambda_{s}\phi_{s}$ so that $(\rho_{q}-\mathrm{Id})\big(\sum_{s}\lambda_{s}\phi_{s}\big)\equiv 0$ on $\mathcal{K}$
$\displaystyle\triangleright$     Verify $(\rho_{j}-\mathrm{Id})$ vanishes for all $j=1,\dots,q$ on $\mathcal{K}$; accept the invariant if passed
$\displaystyle\triangleright$   end for
$\displaystyle\triangleright$  end if
$\displaystyle\triangleright$  [Step 5] Correction from lifts $\psi(g)$ in target (optional library):
$\displaystyle\triangleright$  Note (library scope): the target-invariant library is optional and may include cases such as $(q,n_{\rm tgt})=(6,15)$ alongside any others that are provided
$\displaystyle\triangleright$  for each known $GL(q)$-invariant $g$ in target degree $n_{\rm tgt}$ do
$\displaystyle\triangleright$   Lift by inverse Kameko: $\psi(g)=\sum x_{1}^{2e_{1}+1}\cdots x_{q}^{2e_{q}+1}$ for each monomial $x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}$ in $g$
$\displaystyle\triangleright$   Let $\mathcal{L}:=\{i\in\mathcal{K}:\text{weight}(i)<\text{weight}(\psi(g))\}$
$\displaystyle\triangleright$   Repeat Stage 1/2 on the subset $\mathcal{K}$ for $\psi(g)$, verify $\rho_{j}$-invariance; collect accepted invariants
$\displaystyle\triangleright$  end for
$\displaystyle\triangleright$  Output: union of all accepted $GL(q)$-invariants from Step 4 and Step 5 (with logs of weights and dimensions)
$\displaystyle\trianglerightend$ procedure
Remark 3.3 (Key techniques and why they matter).
  • 1.
    Streaming + ONLINE hit elimination (pivot map). Instead of assembling the full Steenrod action matrix and performing Gaussian elimination, the algorithm streams each column $Sq^{2^{p}}(M)$, maps monomials to row indices, and performs online XOR-reduction against a sparse pivot map. This directly constructs an admissible basis of $QP_{q}$ in degree $n$ with a controlled memory footprint and scales well for large $(q,n)$.
  • 2.
    Ordering by weight vector $\omega$ and weight grouping. Sorting exponent tuples by $(\omega(\mathbf{a}),\mathbf{a})$ yields a canonical admissible basis and enables blockwise decomposition by weight. Subsequent linear systems (for $\Sigma_{q}$ and $GL(q)$) are then solved weight-by-weight, which substantially reduces system sizes.
  • 3.
    Bit-packed matrices and Gaussian elimination over $\mathbb{F}_{2}$. All nullspace computations (Kameko matrix, $\Sigma_{q}$-stage, $GL(q)$-stage) use bit-packed matrices, so elimination and back-substitution become word-level XOR operations. This is cache-friendly and significantly faster than dense arithmetic over $\mathbb{F}_{2}$.
  • 4.
    Kameko map at the level of exponents. The Kameko matrix $L$ is built via exponent arithmetic: $\mathbf{a}\mapsto(\mathbf{a}-\mathbf{1})/2$ when all entries are odd, followed by reduction to admissible rows in the target degree. This avoids heavy polynomial manipulation while preserving the required linear structure.
  • 5.
    Precomputation of $(\rho_{j}-\mathrm{Id})$ rows on the admissible basis. For each $j=1,\dots,q$ and each admissible monomial $u_{i}$, the row support of $(\rho_{j}-\mathrm{Id})u_{i}$ is computed once and reused across the $\Sigma_{q}/GL(q)$ stages and the Stage 4–5 corrections, eliminating repeated decompositions.
  • 6.
    Subset-based correction within the kernel and under weight constraints. In Stage 4–5 the derivation is solved only on the admissible indices that lie in the Kameko kernel support and, when appropriate, only against lower-weight monomials than $w(g)$. This turns global constraints into a few smaller, sparse systems on restricted index sets.
[…]
Now, by applying the above algorithm for $q=6$ and $n=36,$ we obtain an isomorphism:
${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})\cong\bigoplus_{1\leq i\leq 5}QP_{6}(\omega^{*}_{(i)}),$
where
$\displaystyle\omega^{*}_{(1)}:=(4,2,1,1,1),\ \ \omega^{*}_{(2)}:=(4,2,1,3),\ \ \omega^{*}_{(3)}:=(4,2,3,2),$
$\displaystyle\omega^{*}_{(4)}:=(4,4,2,2),\ \ \omega^{*}_{(5)}:=(4,4,4,1).$
Then, our algorithm finds:
$i$ 1 2 3 4 5
$\dim QP_{6}(\omega^{*}_{(i)})$ 2725 111 1085 6495 1974
Thus, $\dim{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})=\sum_{1\leq i\leq 5}\dim QP_{6}(\omega^{*}_{(i)})=12390.$
Using the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 5,$ we get:
$i$ 1 2 3 4 5
$\dim[QP_{6}(\omega^{*}_{(i)})]^{\Sigma_{6}}$ 13 2 6 18 13
Using the homomorphism $\rho_{6}:P_{6}\longrightarrow P_{6},$ we obtain:
$\dim[QP_{6}(\omega_{(i)})]^{GL(6)}=\left\{\begin{array}[]{ll}1&\mbox{if $\displaystyle i=1,\,5$},\\ 0&\mbox{otherwise}.\end{array}\right$.
Furthermore, $[QP_{6}(\omega_{(1)})]^{GL(6)}=\mathbb{F}_{2}\cdot[\zeta_{1}],$ and $[QP_{6}(\omega_{(5)})]^{GL(6)}=\mathbb{F}_{2}\cdot[\widetilde{\zeta}]_{\omega_{(5)}},$ where the polynomial $\zeta_{1}$ is given as in Section 1, and
$\displaystyle\widetilde{\zeta}$ $\displaystyle=x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{14}x_{5}^{3}x_{6}^{5}+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{6}x_{5}^{6}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{6}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{10}x_{5}^{6}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{14}x_{5}^{6}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{6}x_{6}^{5}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{5}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{10}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{14}x_{6}^{5}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{14}x_{6}^{5}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{5}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{6}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{6}x_{5}^{5}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{10}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{10}x_{5}^{5}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{5}x_{6}^{6}+x_{1}^{7}x_{2}^{9}x_{3}^{3}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{13}x_{3}^{3}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{9}x_{4}^{5}x_{5}^{6}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{13}x_{4}^{5}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{9}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{9}x_{5}^{6}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{9}x_{5}^{6}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{13}x_{5}^{6}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{9}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{9}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{9}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{6}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{5}x_{5}^{12}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{13}x_{6}^{6}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{13}x_{6}^{6}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{7}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{12}x_{6}^{7}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{12}x_{6}^{7}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{14}x_{6}^{7}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{14}x_{6}^{7}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{7}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{14}x_{6}^{7}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{6}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{7}x_{5}^{6}x_{6}^{9}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{6}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{6}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{7}x_{5}^{3}x_{6}^{12}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{6}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{7}x_{5}^{6}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{7}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{7}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{6}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{6}x_{6}^{13}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{6}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{5}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{5}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{5}x_{6}^{14}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{7}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{7}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{7}x_{6}^{14}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{7}x_{6}^{14}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{14}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{7}x_{6}^{14}.$
Assume that $g\in P_{6}$ such that $[g]\in[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)},$ then
$g\equiv\gamma\widetilde{\zeta}+\sum_{x\in\mathbf{Ad}_{6}(\omega_{(i)}),\,1\leq i\leq 4}\gamma_{x}\cdot x,\ \ \gamma,\,\gamma_{x}\in\mathbb{F}_{2},$
where $|\mathbf{Ad}_{6}(\omega_{(i)}|=\dim QP_{6}(\omega^{*}_{(i)})$ for all $i,\,1\leq i\leq 4,$ and the set of all admissible monomials in $\mathbf{Ad}_{6}(\omega_{(i)}$ has also been listed in detail in the output of the algorithm as in Note 3.5(B). Using the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 5,$ and the relation $\rho_{j}(g)\equiv g,$ we see that
$g\equiv\gamma(\widetilde{\zeta}+h_{0})+\ \ \mbox{$\displaystyle 517$ terms $\displaystyle\beta_{i}g^{\prime\prime}_{i}$},\ \ \gamma,\,\beta_{i}\in\mathbb{F}_{2},$
where the polynomials $g^{\prime\prime}_{i},\,1\leq i\leq 517,$ are determined from the algorithm output in Note 3.5(B), and
$\displaystyle h_{0}$ $\displaystyle=x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{9}x_{5}^{9}x_{6}^{8}+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{7}x_{4}^{3}x_{5}^{10}x_{6}^{8}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{8}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{3}x_{5}^{12}x_{6}^{8}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{10}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{6}x_{6}^{9}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{9}x_{5}^{8}x_{6}^{9}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{8}x_{5}^{9}x_{6}^{9}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{8}x_{3}^{3}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{14}x_{3}^{3}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{8}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{8}x_{4}^{5}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{7}x_{3}^{10}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{11}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{6}x_{3}^{11}x_{4}^{5}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{14}x_{4}^{5}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{8}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{8}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{12}x_{5}^{3}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{13}x_{5}^{3}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{14}x_{5}^{3}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{14}x_{5}^{3}x_{6}^{10}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{8}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{8}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{10}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{11}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{11}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{14}x_{5}^{5}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{9}x_{5}^{6}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{11}x_{5}^{6}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{11}x_{5}^{6}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}x_{4}^{13}x_{5}^{6}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{7}x_{3}^{5}x_{4}^{3}x_{5}^{8}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{7}x_{4}^{4}x_{5}^{9}x_{6}^{10}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{6}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{7}x_{5}^{9}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{11}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{11}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{10}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{11}x_{6}^{10}+x_{1}^{7}x_{2}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{10}+x_{1}x_{2}^{7}x_{3}^{3}x_{4}^{3}x_{5}^{12}x_{6}^{10}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{12}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{5}x_{5}^{12}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{13}x_{6}^{10}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{13}x_{6}^{10}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{13}x_{6}^{10}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}^{3}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{6}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{8}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{6}x_{4}x_{5}^{10}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{4}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{4}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}x_{2}^{6}x_{3}^{3}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{5}x_{5}^{10}x_{6}^{11}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{5}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{6}x_{5}^{10}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{11}+x_{1}x_{2}^{3}x_{3}^{5}x_{4}^{6}x_{5}^{10}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{11}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{12}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}x_{3}^{6}x_{4}x_{5}^{14}x_{6}^{11}+x_{1}^{3}x_{2}^{5}x_{3}x_{4}^{2}x_{5}^{14}x_{6}^{11}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}x_{4}^{6}x_{5}^{14}x_{6}^{11}+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{8}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{8}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{10}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{11}x_{5}^{3}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{11}x_{5}^{3}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{14}x_{5}^{3}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{14}x_{5}^{3}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{7}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{7}x_{3}^{3}x_{4}^{3}x_{5}^{8}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{5}x_{4}^{3}x_{5}^{10}x_{6}^{12}+x_{1}^{3}x_{2}x_{3}^{3}x_{4}^{7}x_{5}^{10}x_{6}^{12}$
$\displaystyle\quad+x_{1}x_{2}^{3}x_{3}^{3}x_{4}^{7}x_{5}^{10}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}^{3}x_{3}^{4}x_{4}^{3}x_{5}^{11}x_{6}^{12}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{11}x_{6}^{12}+x_{1}^{3}x_{2}^{5}x_{3}^{3}x_{4}^{2}x_{5}^{10}x_{6}^{13}+x_{1}^{3}x_{2}^{5}x_{3}^{2}x_{4}^{3}x_{5}^{10}x_{6}^{13}$
$\displaystyle\quad+x_{1}^{3}x_{2}x_{3}^{6}x_{4}^{3}x_{5}^{10}x_{6}^{13}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{10}x_{6}^{13}+x_{1}x_{2}x_{3}^{3}x_{4}^{6}x_{5}^{12}x_{6}^{13}+x_{1}x_{2}^{3}x_{3}^{6}x_{4}^{3}x_{5}^{9}x_{6}^{14}.$
Using the relation $\rho_{6}(g)\equiv g$, we impose the final condition for $[g]$ to be $GL(6)$-invariant. This leads to a system of linear equations over $\mathbb{F}_{2}$ for the coefficients $(\gamma,\beta_{1},\dots,\beta_{517})$ that define the $\Sigma_{6}$-invariant elements. By solving this system, our algorithm finds that the solution space for the coefficients is two-dimensional, and
$g\equiv c_{1}\zeta_{1}+c_{2}\zeta_{2},\ \ \mbox{for some scalars $\displaystyle c_{1},c_{2}\in\mathbb{F}_{2}$},$
where the polynomials $\zeta_{1}$ and $\zeta_{2}$ are determined as in Section 1. Thus, the calculations show that
$\dim[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}=2,$
and
${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}=\mathbb{F}_{2}\cdot([\zeta_{1}],[\zeta_{2}]).$
Remark 3.4.
Since ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ is a subspace of $(QP_{6})_{36}$ and the non-zero elements $[\zeta_{1}]$ and $[\zeta_{2}]$ are $GL(6)$-invariants in ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)}),$ they are also $GL(6)$-invariants in $(QP_{6})_{36}.$ This implies that
$\dim(\mathbb{F}_{2}\otimes_{GL(6)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{6})))_{36}=\dim[(QP_{6})_{36}]^{GL(6)}>\dim{\rm Ext}^{6,6+36}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})=1.$
Furthermore, our algorithm also finds that $[(QP_{6})_{36}]^{GL(6)}=\mathbb{F}_{2}\cdot([\zeta_{1}],[\zeta_{2}]).$ Indeed, using this result and Proposition 3.2, we see that if $h\in P_{6}$ such that $[h]\in[(QP_{6})_{36}]^{GL(6)},$ then
$h\equiv\beta\psi(\xi)+h^{*},\ \ \beta\in\mathbb{F}_{2},$
where the polynomial $\xi$ is determined as in Proposition 3.2, $\psi$ is the Kameko lift homomorphism $(P_{6})_{15}\longrightarrow(P_{6})_{36},\,x_{1}^{e_{1}}\ldots x_{6}^{e_{6}}\longmapsto x_{1}^{2e_{1}+1}\ldots x_{6}^{2e_{6}+1},$ and $h^{*}\in P_{6}$ such that $[h^{*}]\in\mathrm{Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)}).$ Then our algorithm finds that $\beta=0$ and $h\equiv d_{1}\zeta_{1}+d_{2}\zeta_{2},$ for some scalars $d_{1},d_{2}\in\mathbb{F}_{2}.$ Therefore,
$\dim[(QP_{6})_{36}]^{GL(6)}=\dim[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}=2,$
and
$(QP_{6})_{36}]^{GL(6)}=\mathbb{F}_{2}\cdot([\zeta_{1}],[\zeta_{2}]).$
By direct manual verification with computer assistance, we also obtain $\rho_{i}(\zeta_{1})\equiv\zeta_{1}$ and $\rho_{i}(\zeta_{2})\equiv\zeta_{2}$ for all $i,\,1\leq i\leq 6.$ This completes the proof of the theorem.
Note 3.5.
We have also conducted cross-validation of the results computed manually in our previous work, and our algorithm yields output that demonstrates complete consistency with those results. The explicit computational code implemented in OSCAR is available upon request.
  1. (A)
    The detailed output for the case $q=6$, $n=15$ is available at:
    https://doi.org/10.5281/zenodo.17620661.
  2. (B)
    The detailed output for the case $q=6$, $n=36$ is available at:
    https://doi.org/10.5281/zenodo.17620680.
  3. (C)
    Why OSCAR instead of SageMath [29]? We chose to implement the present algorithm in OSCAR (built on Julia, Nemo/AbstractAlgebra, and FLINT) rather than in SageMath, for the following technical reasons that are directly aligned with our workload:
    • (a)
      Just-in-time compiled inner loops. The streaming hit-elimination, bit-packed Gaussian elimination over $\mathbb{F}_{2}$, and weight-wise kernels are implemented as type-stable Julia loops. This avoids the interpreter overhead of pure Python-level iterations and allows the compiler to inline and vectorize critical sections.
    • (b)
      Bit-level linear algebra. Our nullspace routine operates on packed UInt64 rows with branch-free XOR sweeps. Julia’s low-level bit operations map cleanly to machine code, yielding high throughput for large, very sparse $\mathbb{F}_{2}$ systems.
    • (c)
      Thread-parallel sections. Where safe (e.g. independent column builds, precomputation of $(\rho_{j}-\mathrm{Id})$ rows), we use Base.Threads to parallelize without introducing global-interpreter locks. This is effective for the combinatorial enumeration that dominates running time.
    • (d)
      Tight integration with polynomial arithmetic over $\mathbb{F}_{2}$. Via OSCAR/Nemo, monomial and polynomial operations (Kameko images, Steenrod squares with Lucas’ criteria) are executed by libraries optimized in C/Julia, reducing allocation and dispatch overhead.
    • (e)
      Memory-aware streaming. The ONLINE elimination uses adaptive batching driven by live-heap estimates (soft/hard thresholds), so large degrees can be processed without constructing dense matrices in memory. This design is natural to express in Julia and integrates well with the GC (Garbage Collector) and logging.
    We emphasize that the mathematical pipeline is platform-agnostic: the streaming hit elimination, Kameko kernel, weightwise $\Sigma_{q}$$GL(q)$ analysis, and the two-stage corrections (Steps 4-5) can be reproduced in SageMath. In our experience, however, the combination of compiled inner loops, bit-packed algebra, and thread-parallel precomputations in OSCAR leads to markedly faster and more memory-stable runs on the large instances considered here.
    In particular, we construct an algorithm that computes the $GL(q)$-invariants of $(QP_{q})_{n}$ for arbitrary $q$ and $n$ independently of the usual route via the Kameko homomorphism (i.e., without computing invariants of its kernel). This algorithm was initially implemented in SageMath [29] and has since been ported to OSCAR; the source code is available upon request.
Funding
D- ặng Võ Phúc was funded by the Post-Doctoral Scholarship Programme of Vingroup Innovation Foundation (VINIF), Institute of Big Data, code: VINIF.2024.STS.38.
Declaration of Competing Interests
The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The data supporting the findings of this study (specifically, the detailed computational outputs for the cases $(q=5,n=35),$ $(q=6,n=15),$ and $(q=6,n=36)$) are openly available at the URLs provided in the manuscript. The OSCAR source code developed for the computations is available from the corresponding author upon reasonable request.